# DeFlow: Decoupling Manifold Modeling and Value Maximization for Offline Policy Extraction

### v1

@@ 3 DeFlow: Disentangling Fidelity and Optimality
   $a^{\prime}\sim$ sample_action($s$)
@@
   Update $\psi$ to minimize $\mathbb{E}[\|\mu_{\color[rgb]{0.9609375,0.2578125,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.2578125,0.328125}\psi}(t,s,x^{t})-(x^{1}-x^{0})\|_{2}^{2}]$
@@
   Update $\phi$ to minimize $\mathbb{E}[-Q_{\phi}(s,a_{\text{base}}+\Delta)+\alpha\|\Delta\|_{2}^{2}]$
@@ 5 Experiments
Table 1: Offline RL results. DeFlow achieves state-of-the-art or competitive performance across $\mathbf{73}$ diverse benchmark tasks, significantly outperforming baselines on complex environments such as cube-double-play and puzzle-3x3-play. We follow the standard evaluation protocol from fql_park2025. Results are averaged over 8 seeds (4 for pixel-based tasks) and the values are taken from prior works (corl_tarasov2023; idql_hansenestruch2023; srpo_chen2024; fql_park2025). See Table˜6 for the full results.
@@ 5.1 Experimental Setup
relocate-cloned-v1 0 $\to$ 0 11 $\to$ 77 0-0 $\to$ 0-0 0 $\pm 0$ $\to$ 22 $\pm 2$ 0-0 $\pm 0$ $\to$ 55 $\pm 3$ 0 $\pm 1$ $\to$ $\mathbf{62}$ $\pm 8$ $\mathbf{1}$ $\pm 0$ $35$ $\pm 8$
@@ 5.2 Results and Analysis
Yes, by treating it as a semantic constraint rather than an abstract weight. Unlike the behavioral cloning weight $\alpha$, which often requires extensive tuning, $\delta$ represents a physical constraint reflecting the action manifold width. To quantify this, we compute the Intrinsic Action Variance (IAV, denoted as $\bar{\sigma}^{2}_{\text{data}}$) using the average variance of actions among $k$-nearest neighbors. Empirically, we set $\delta$ proportional to the estimated deviation (e.g., $\delta\approx 0.1\bar{\sigma}_{\text{data}}$ for fine manipulation and $\approx\bar{\sigma}_{\text{data}}$ for navigation). Crucially, our adaptive Lagrangian multiplier ensures robustness to rough estimates of $\delta$, effectively eliminating the need for grid search. [^1] For standard D4RL tasks, $\delta$ can simply be set based on data quality, e.g., $10^{-3}$ for expert datasets. See Section A.1 for details.
@@
A: Table˜2 and Figure˜3 summarize the results across $15$ tasks. Quantitatively, DeFlow matches or exceeds the state-of-the-art (e.g., FQL), achieving remarkable gains in complex tasks like humanoidmaze-medium. A critical advantage of DeFlow is its seamless transition from offline to online. Unlike prior works (fql_park2025; agrawalla2025floq) that often necessitate altering hyperparameters (or “loosening” constraints) to encourage exploration, DeFlow requires zero additional tuning. We maintain the exact offline configuration, relying on our automatic $\alpha$ tuning to naturally balance exploration. While this strict adherence to a unified protocol may limit performance in tasks with poor data quality (e.g., relocate-cloned), it ensures empirical rigor and demonstrates the method’s inherent robustness without task-specific engineering.
@@
A: Yes. Our decoupled design offers a unique efficiency benefit: we can freeze the generative behavior prior (Base Flow) and solely rely on the adaptive $\alpha$ to modulate exploration. As shown in Figure˜3, freezing the prior yields performance competitive with full fine-tuning. This strategy reduces computational overhead (runtime analysis provided in Appendix˜B) and enhances stability by eliminating the moving target problem. Crucially, it provides a scalable pathway for fine-tuning large-scale foundation models, where updating the generative backbone online is computationally prohibitive.
@@
A: Yes. While our focus is on policy extraction, DeFlow’s refinement framework is orthogonal to value modeling techniques like floq (agrawalla2025floq). As shown in Table˜3, we directly integrate DeFlow with floq’s value function and observe synergistic improvements without any tuning, suggesting that DeFlow can serve as a versatile policy extraction module compatible with advanced value estimators.
@@ A.1 Hyperparameter Settings
This heuristic contrasts sharply with the exhaustive grid search required for the behavioral cloning weight $\alpha$ in prior works (fql_park2025). Our method relies only on rough estimates of $\delta$ inferred from task semantics, drastically reducing tuning effort. Crucially, we maintain the same $\delta$ for online fine-tuning without further adjustment. While this strict protocol may limit asymptotic performance on specific outlier tasks, it underscores the method’s inherent robustness and eliminates the need for task-specific engineering during the offline-to-online transition. Detailed hyperparameter settings for all experiments are provided in Table˜5.
@@ Appendix B Additional Experimental Results
We report the full offline RL results across all $73$ tasks in Table˜6. The results are averaged over 8 seeds (4 seeds for pixel-based tasks) and we bold scores that are within 95% of the best performance as in FQL (fql_park2025).
Table˜4 compares the training and inference speeds of FQL and DeFlow. The run times are measured on the same machine with an NVIDIA 3090 GPU. Because DeFlow requires the multi-step flow to take samples for Q updates during training, its training time is slightly longer than FQL (while their actor update times are the same). During inference, when the multi-step flow is fixed and only the refinement module is used, DeFlow’s inference time can be faster, which will be more significant when scaling up the flow model.

### v2

@@ 3 DeFlow: Disentangling Fidelity and Optimality
   $a^{\prime}\sim$ sample_action($s^{\prime}$)
@@
   Update $\psi$ to minimize $\mathbb{E}[\|v_{\color[rgb]{0.9609375,0.2578125,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.2578125,0.328125}\psi}(t,s,x^{t})-(x^{1}-x^{0})\|_{2}^{2}]$
@@
   Update $\phi$ to minimize $\mathbb{E}[-Q^{\text{Normalized}}_{\phi}(s,a_{\text{base}}+\Delta)+\alpha\|\Delta\|_{2}^{2}]$
@@ 5 Experiments
Table 1: Offline RL results. DeFlow achieves state-of-the-art or competitive performance across $\mathbf{73}$ diverse benchmark tasks, significantly outperforming baselines on complex environments such as cube-double-play and puzzle-3x3-play. We follow the standard evaluation protocol from fql_park2025. Results are averaged over 8 seeds (4 for pixel-based tasks) and the values are taken from prior works (corl_tarasov2023; idql_hansenestruch2023; srpo_chen2024; fql_park2025). See Table˜7 for the full results.
@@ 5.1 Experimental Setup
relocate-cloned-v1 0 $\to$ 0 11 $\to$ 77 0-0 $\to$ 0-0 0 $\pm 0$ $\to$ 22 $\pm 2$ 0-0 $\pm 0$ $\to$ 55 $\pm 3$ 0 $\pm 1$ $\to$ $\mathbf{62}$ $\pm 8$ $\mathbf{1}$ $\pm 0$ $\to$ $35$ $\pm 8$
@@ 5.2 Results and Analysis
Yes, by treating it as a semantic constraint rather than an abstract weight. Unlike the behavioral cloning weight $\alpha$, which often requires extensive tuning, $\delta$ represents a physical constraint reflecting the action manifold width. To quantify this, we compute the Intrinsic Action Variance (IAV, denoted as $\bar{\sigma}^{2}_{\text{data}}$) using the average variance of actions among $k$-nearest neighbors. Empirically, we set $\delta$ proportional to the estimated deviation (e.g., $\delta\approx 0.1\bar{\sigma}_{\text{data}}$ for fine manipulation and $\approx\bar{\sigma}_{\text{data}}^{2}$ for navigation). Crucially, our adaptive Lagrangian multiplier ensures robustness to rough estimates of $\delta$, effectively eliminating the need for grid search. [^1] For standard D4RL tasks, $\delta$ can simply be set based on data quality, e.g., $10^{-3}$ for expert datasets. See Section A.1 for details.
@@
A: Table˜2 and Figure˜3 summarize the results across $15$ tasks. Quantitatively, DeFlow matches or exceeds the state-of-the-art (e.g., FQL), achieving remarkable gains in complex tasks like humanoidmaze-medium. A critical advantage of DeFlow is its seamless transition from offline to online. Unlike prior works (fql_park2025; agrawalla2025floq) that often necessitate altering hyperparameters (or “loosening” constraints) to encourage exploration, DeFlow requires zero additional tuning. We maintain the exact offline configuration, relying on our automatic $\alpha$ tuning to naturally balance exploration. While this strict adherence to a unified protocol may limit performance in tasks with poor data quality (e.g., relocate-cloned), it ensures empirical rigor and demonstrates the method’s inherent robustness without task-specific engineering. Additional ablation studies are provided in Section˜B.2.
@@
A: Yes. Our decoupled design offers a unique efficiency benefit: we can freeze the generative behavior prior (Base Flow) and solely rely on the adaptive $\alpha$ to modulate exploration. As shown in Figure˜3, freezing the prior yields performance competitive with full fine-tuning. This strategy reduces computational overhead (runtime analysis provided in Section˜B.1) and enhances stability by eliminating the moving target problem. Crucially, it provides a scalable pathway for fine-tuning large-scale foundation models, where updating the generative backbone online is computationally prohibitive.
@@
A: Yes. While our focus is on policy extraction, DeFlow’s refinement framework is orthogonal to value modeling techniques like floq (agrawalla2025floq). As shown in Table˜3, we integrate DeFlow with floq’s value function and observe synergistic improvements without any tuning, suggesting that DeFlow can serve as a versatile policy extraction module compatible with advanced value estimators.
@@ A.1 Hyperparameter Settings
This heuristic contrasts sharply with the exhaustive grid search required for the behavioral cloning weight $\alpha$ in prior works (fql_park2025). Our method relies only on rough estimates of $\delta$ inferred from task semantics, drastically reducing tuning effort. Crucially, we maintain the same $\delta$ for online fine-tuning without further adjustment. While this strict protocol may limit asymptotic performance on specific outlier tasks, it underscores the method’s inherent robustness and eliminates the need for task-specific engineering during the offline-to-online transition. Detailed hyperparameter settings for all experiments are provided in Table˜6.
@@ Appendix B Additional Experimental Results
We report the full offline RL results across all $73$ tasks in Table˜7. The results are averaged over 8 seeds (4 seeds for pixel-based tasks) and we bold scores that are within 95% of the best performance as in FQL (fql_park2025).
B.1 Runtime Analysis
Table˜4 compares the training and inference speeds of FQL and DeFlow. The run times are measured on the same machine with an NVIDIA 3090 GPU. Because DeFlow requires the multi-step flow to take samples for Q updates during training, its training time is slightly longer than FQL (while their actor update times are the same). During online training, when the multi-step flow is fixed and only the refinement module is used, DeFlow’s training time can be faster, which will be more significant when scaling up the flow model.
@@ B.1 Runtime Analysis
B.2 Ablations on $\alpha$ Adaptation
To validate the effectiveness of our adaptive Lagrangian multiplier $\alpha$, we conduct ablation studies on cube-double-play and puzzle-3x3-play tasks by comparing with fixed $\alpha$ values.
Table 5: Ablation study on fixed vs. adaptive $\alpha$ values.
Tasks $\alpha=0.03$ $0.1$ $0.3$ 11 33 Adaptive
cube-double-play-singletask-v0 0 $\pm 0$ $\to$ 0 $\pm 0$ 0 $\pm 0$ $\to$ 0 $\pm 0$ 0 $\pm 0$ $\to$ 0 $\pm 0$ $44$ $\pm 2$ $\to$ $10$ $\pm 3$ 77 $\pm 2$ $\to$ $89$ $\pm 5$ $\mathbf{48}$ $\pm 13$ $\to$ $\mathbf{93}$ $\pm 4$
puzzle-3x3-play-singletask-v0 0 $\pm 0$ $\to$ $\mathbf{100}$ $\pm 1$ 0 $\pm 0$ $\to$ $\mathbf{100}$ $\pm 0$ 77 $\pm 7$ $\to$ $\mathbf{100}$ $\pm 0$ 0 $\pm 0$ $\to$ 0 $\pm 0$ $\mathbf{35}$ $\pm 8$ $\to$ 0 $\pm 0$ $23$ $\pm 8$ $\to$ $\mathbf{100}$ $\pm 0$
We observe a significant trade-off when using fixed $\alpha$. Low $\alpha$ values (e.g., $\alpha<1$) fail to enforce the constraint in precise tasks like cube-double-play, leading to severe OOD issues and near-zero performance ($0\to 0$). Conversely, overly high $\alpha$ values (e.g., $\alpha=3$) impose excessive regularization; while this ensures offline stability in puzzle-3x3-play ($35$), it hinders online exploration, causing performance to collapse to 0. In contrast, our Adaptive mechanism dynamically adjusts $\alpha$ to balance constraint satisfaction and reward maximization. It achieves superior or comparable performance across both tasks without the need for task-specific tuning.